The Evanec Company's next expected dividend, D1, is $2.92; its growth rate is 4%; and its common stock now sells for $37.00. New stock (external equity) can be sold to net $33.30 per share.

  1. What is Evanec's cost of retained earnings, rs? Do not round intermediate calculations. Round your answer to two decimal places.

    rs =   %

  2. What is Evanec's percentage flotation cost, F? Round your answer to two decimal places.

    F =   %

  3. What is Evanec's cost of new common stock, re? Do not round intermediate calculations. Round your answer to two decimal places.

    re =   %
Answer #1

a.Cost of retained earnings=(D1/Current price)+Growth rate

=(2.92/37)+0.04

=11.89%(Approx).

b.% flotation cost=(Current price-Net price)/Current price

=(37-33.3)/37

=10%

c.Cost of new common stock=(D1/Net price)+Growth rate

=(2.92/33.3)+0.04

=12.77%(Approx).